# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ V Results II: Power-law divergence of correlation length at high temperature
As the system approaches the critical point of a BKT transition from a disordered phase, $\xi$ and $\chi$ diverge exponentially as $\sim\exp(b/\sqrt{t})$ [49, 50], where $t=(T-T_{c})/T_{c}$ is the reduced temperature, and $b$ is a non-universal constant. As shown in Fig. 10(a), a semi-logarithmic plot of $\xi$ versus $a+b\sqrt{t}$ is presented for $\sigma=2.1$ (purple), $3$ (red), and the NN case (black), where the non-universal constant $a$ comes from the amplitude of the scaling $\sim\exp(b/\sqrt{t})$, and the specific values of $a$ and $b$ are provided in the corresponding caption.
In the critical scaling regime, for a given $t$, $\xi$ quickly converges to its thermodynamical value as $L$ increases. The growth of $\xi$ at large $L$ as a function of $t$ then reveals the thermodynamic growth law of correlation length; however, when the system enters the finite-size critical window, where $\xi\sim L$, the $\xi$ curve begins to bend to a plateau due to finite-size scaling behavior, thus deviating from the thermodynamic growth law. Note that this analysis is different from the FSS analysis. Here, one first takes the $L\rightarrow\infty$ limit for a given temperature $T$ and then takes the $T\rightarrow T_{c}$ limit to examine the asymptotic growth law of $\xi$. In contrast, the FSS analysis first takes the $T\rightarrow T_{c}$ limit for a given $L$, followed by the $L\rightarrow\infty$ limit, studying the scaling behavior within the finite-size critical window. As shown in Fig. 10(a), the straight black line in the semi-log plot clearly demonstrates the exponential growth of $\xi$ for $\sigma>2$, the typical BKT scaling behavior. In the large $t$ regime (the bottom left corner of Fig. 10(a)), where the system is far from criticality, $\xi$ slightly deviates from the exponential growth. Moreover, in Fig. 10(b), we further plot the ratio $\xi/L$ versus $b/\sqrt{t_{L}}=b/\sqrt{t(\ln L/L_{0})^{2}}$ to examine the FSS of the expoential growth, where $L_{0}$ is non-universal length scale, simply set as $1$. The scaling field $t_{L}$ [51, 52] originates from the exponential divergence of the correlation length near the BKT transition $\xi\sim L\sim\exp(b/\sqrt{t})$. The data points collapse onto a single curve, showing that for $\sigma>2$, the phase transitions belong to the BKT universality class.
@@
In Fig. 11(a), a semi-logarithmic plot of $a\xi$ versus $b/\sqrt{t}$ for various $\sigma$ is presented. The non-universal constants $a$ and $b$, obtained through the fitting, enhance the overall readability of plots without affecting the universal growth law. The specific values of $a$ and $b$ are provided in the corresponding caption. For $\sigma=2.1$ (purple), $\sigma=3$ (red), and the NN case (black), the straight black line demonstrates that the correlation length diverges exponentially $\xi\sim\exp(b/\sqrt{t})$, characterizing a typical BKT transition. For clarity, we exclude certain data points in the finite-size critical window, which manifests as a plateau. For $\sigma=1.875$ (green) and $2$ (blue), when sufficiently away from $\beta_{c}$, i.e. $b/\sqrt{t}$ approaches 0, $\xi$ behaves seemingly like that for $\sigma=2.1$, $3$ and the NN case. However, near the critical point, i.e., as $b/\sqrt{t}$ approaches infinity, the behavior of $\xi$ becomes increasingly different from SR cases ($\sigma>2$) and deviates from the exponential growth, suggesting a different universality class. In fact, the growth of $\xi$ for $\sigma=1.875$ and $2$ can be well-fitted using a power-law function (the curved lines in dark green and dark blue), indicating a second-order transition. In Fig. 11(b), a similar analysis is applied to the susceptibility, $\chi$. The plot shows $a\chi$ versus $b/\sqrt{t}$ on a semi-logarithmic scale. For the BKT cases ($\sigma=2.1$, $3$, and NN), $\chi$ diverges exponentially, with the straight black line guiding the eye. On the other hand, for $\sigma=1.875$ and $2$, $\chi$ follows a power-law growth near the critical point, with the dark-blue and dark-green curved lines guiding the eye, suggesting a second-order phase transition.
Figure. 12 presents the log-log plot of $a\xi$ and $a\chi$ versus ${b^{2}}/{t}$, where the x-axis is the square of ${b}/{\sqrt{t}}$. In Fig. 12(a), for $\sigma=1.875$ and $2$, the growth of $\xi$ asymptotically follows a linear relation as ${b^{2}}/{t}\rightarrow\infty$ (the straight dark-green and dark-blue lines), consistent with the power-law growth behavior of a second-order transition. The deviation of $\xi$ from linearity and the discrepancy between the exponent estimated from Fig. 12(a) and fitting result in Table 5 are likely due to the strong correction to scaling from the irrelevant fields. Specifically, $\xi\sim t^{-\nu}g(u/t^{y_{u}/y_{t}})$, where $u$ is some irrelevant field and $y_{u}<0$ is the correpsonding scaling exponent. Nevertheless, the growth behavior of $\xi$ can clearly reveal the type of phase transition for various $\sigma$. Similarly, for the growth of $\chi$ in Fig. 12(b), the straight dark-green and dark-blue lines illustrate a power-law behavior, as $\chi\sim t^{-\gamma}$ for $\sigma=1.875$ and $2$, again indicating a second-order transition. As ${b^{2}}/{t}$ approaches infinity, the slopes provide the critical exponent $\gamma$, and through the hyperscaling relation $\gamma=\nu(2-\eta)$, these values are also consistent with those listed in Table 5. In all, our analysis of the growth behavior of $\xi$ and $\chi$ strongly suggests that the system undergoes a second-order transition for $\sigma\leq 2$, while a BKT transition for $\sigma>2$; hence, the threshold between SR and LR universality is at $\sigma_{*}=2$.
VI Results III: Critical points and critical exponents
@@
Figure 13: The zoom-in of the 2nd-moment correlation length divided by $L$, $\xi/L$, as a function of inverse temperature $\beta$ for various system sizes, ranging from $L=64$ to $L=4096$, is shown across three different values of $\sigma$: $1.75$ (a), $1.875$ (b), and $2$ (c). Clear crossing behavior can be observed in all the cases, despite the shift of crossing point due to the finite-size effect.
@@
Figure 14: The extrapolation of $\xi/L$ curves’ crossing point $\beta_{L}$ of different system sizes. The $x$-axis represents the correction term $L^{-\omega}$: $L^{-0.86}$ in panel (a), $L^{-0.84}$ in panel (b), and $L^{-0.75}$ in panel (c), while the $y$-axis denotes the inverse temperature $\beta$ at which the correlation length ratios for two consecutive system sizes intersect. The red line indicates the extrapolated value of $\beta$ as the system size approaches infinity.
@@
In the previous sections, we have studied the low-$T$ and high-$T$ properties of the system for $\sigma\leq 2$. In this section, we focus on the system’s critical properties for $\sigma\leq 2$. The fitting of the critical point and critical exponent is elaborated in Sec. VI.1 and VI.2, respectively. Then, to determine the reliability of the estimate of $\eta$, the correlation function is also studied in Sec. VI.3.
VI.1 The Fitting of Critical Point
@@
In this subsection, we focus on determining the critical temperature for various values of $\sigma$. For $\sigma\leq 2$, a second-order phase transition from a disordered phase to a ferromagnetic phase is expected, where the dimensionless ratio curves, such as $Q_{m}$ and $\xi/L$, for different system sizes, should ideally intersect at the critical point. Figure 13 presents $\xi/L$ as a function of inverse temperature $\beta$ for system sizes ranging from $L=64$ to $L=4096$, and for $\sigma=1.75$, $1.875$, and $2.0$. Unlike Figure 2, which covers a broader range, this figure zooms in on a small regime of $\beta$ to observe the crossing behavior of $\xi/L$ for different system sizes. However, Figure 13 shows that the crossing points for different system sizes do not converge at a single value of $\beta$. Instead, these points shift toward higher values of $\beta$ as the system size increases, a phenomenon attributed to finite-size effects. The distance between the crossing points decreases as the system size grows, indicating that with larger system sizes, the crossing points will eventually converge at a single value of $\beta$, marking the critical point. To confirm this behavior and accurately determine the critical point, a finite-size analysis is performed to the crossing point $\beta_{L}$ between system sizes of $L$ and $L/2$.
@@
Figure 15: Demonstration of the reliability of the fitting of $\chi$. The semi-logarithmic plots of $\chi L^{\eta-2}-bL^{-\omega}$ versus $L$ for $\sigma=1.75,1.875,$ and $2$ are shown (where $\omega=0.48,0.42,0.5$ respectively). The red square approaches a horizontal line as $L\rightarrow\infty$, indicating the precise estimate for $\eta$ that we obtain.
@@
Figure 16: Demonstration of the reliability of the fitting of $K$. The semi-logarithmic plots of $KL^{-y_{t}}-bL^{-\omega}$ versus $L$ for $\sigma=1.75,1.875,$ and $2$ are shown (where $\omega=0.25,0.25,0.2$ respectively). The red square approaches a horizontal line as $L\rightarrow\infty$, indicating the precise estimate for $y_{t}$ that we obtain (note that $y_{t}=1/\nu$). It is worth noting that, compared to $\chi$, the covariance $K$ exhibits larger errors at various data points. This is particularly evident for $L=8192$, where the difficulty in simulating longer Markov chain lengths results in significant errors for $K$ under this parameter. Therefore, the data point for $L=8192$ is not included in the fitting of $K$.
@@
where $\omega$ is related to the correlation length exponent $\nu$ and the leading irrelevant exponent in the renormalization group framework. Figure 14 illustrates the fitting results, with the $x$-axis representing $L^{-\omega}$ and the $y$-axis showing the corresponding values of $\beta_{L}$. In each case, the fitting lines exhibit finite intercepts on the $y$-axis, highlighted by red stripes. This confirms that the $\xi/L$ curves finally intersect at a single critical point as the system size increases. The intercepts provide an estimate of the critical temperature $\beta_{c}$. For further accuracy, the dimensionless ratios, such as $\xi/L$ and $Q_{m}$, are also fitted with respect to $L$ and $\epsilon=1/T-1/T_{c}$ near the critical point to refine the estimate of the critical point. $\epsilon=\frac{1}{T}-\frac{1}{T_{c}}$ is used to replace $t$ for the convenience of fitting:
$t=a_{1}(T-T_{c})+O((T-T_{c})^{2})=b_{1}\epsilon+O(\epsilon^{2}).$ (39)
Considering the universal part of dimensionless ratio $Q$ as in Eq. (17) and (18), Certain Taylor expansion provides a fitting equation of $Q$:
$Q(t,u,L)$ $=a_{0}+\sum_{i=1}^{m}a_{i}(\epsilon L^{y_{t}})^{i}+b_{1}L^{-\omega}+b_{2}L^{-2\omega}$ (40)
@@
where the value of $m$ depends on the specific fitting process and is typically taken as $2$ or $3$; $L^{-\omega}$ term comes from the irrelevant field $u$, and $\omega=-y_{u}$. In the formula above, terms such as $\epsilon^{2}L^{y_{t}},\epsilon^{3}L^{y_{t}},\epsilon^{3}L^{2y_{t}},\cdots$, are not under consideration because according to experiments, including such terms in the specific fitting process has almost no impact on the fit, and their amplitudes are also very close to zero. For the fit of $Q$, when $\sigma\leq 1.5$, $Q_{m}$ is more suitable for fitting, whereas when $1.5<\sigma\leq 2$, $\xi/L$ provides a better fit. By combining the two quantities, relatively accurate estimates of critical points can be obtained, and they are consistent with those derived from the fit of $\beta_{L}$. The final estimated values of $\beta_{c}$ are presented in Table 4 (Note that, except for $\sigma=1.25,1.5,1.75,1.875$ and $2$, the critical points under other parameters are only determined with much less precision since they are mainly used for drawing the phase diagram in Fig. 1).
Table 4: Estimations of critical point $\beta_{c}$ for various $\sigma$ shown in Fig. 1.
@@
As a supplement, we briefly introduce the method to determine the critical point for $\sigma>2$. When $\sigma>2$, the system undergoes a BKT transition, where the susceptibility $\chi$ is known to follow the scaling behavior: $\chi\sim L^{7/4}(\ln L+C_{1})^{1/8}$ at the critical point [53] where $C_{1}$ is a nonuniversal constant and needs to be obtained through fitting. Therefore, plotting $\chi L^{-7/4}(\ln L+C_{1})^{-1/8}$ as a function of temperature, the curves of different system sizes intersect at the critical point.
VI.2 The Fitting of Critical Exponents
According to the scaling relationship, only two critical exponents are independent, and here we mainly focus on $\eta$ and the thermal scaling exponent $y_{t}$ (note the relationship between $y_{t}$ and the critical exponent $\nu$: $\nu=1/y_{t}$). In Sak’s criterion, $\eta=\max(2-\sigma,\eta_{\rm{SR}})$ where $\eta_{\rm{SR}}=\frac{1}{4}$ for 2D XY model, so estimating the value of $\eta$ is another way to determine whether $\sigma_{*}$ equals 1.75 or 2. By fitting the susceptibility $\chi$ and the scaled covariance $K$, the values of $\eta$ and $y_{t}$ can be numerically estimated. At the critical point, considering the finite-size scaling Eq. (14) and (16), $\chi$ and $K$ are respectively fitted to:
$\chi=L^{2-\eta}(a+b\cdot L^{-\omega})+c$ (41)

### v2

@@ V Power-law divergence of correlation length at high temperature
Figure 14: The extrapolation results of $\xi$ along with the original data for $\sigma=2,3$. The blue dots represent the original data obtained directly from the simulation without extrapolation, as already shown in Fig. 12 and 13; the red dots represent the extrapolated value. Panel (a) is plotted in a double-log coordinate. The distinct linearity of the extrapolated data for $\sigma=2$ demonstrates the power-law divergence of $\xi$: $\xi\sim t^{-\nu}$ with $\nu\approx 2.24$. Such power-law divergence of $\xi$ is the signature of a second-order phase transition. For comparison, in panel (b), $\xi$ is plotted versus $1/\sqrt{t}$ in a semi-log coordinate. The linearity of the extrapolated data for $\sigma=3$ demonstrates the exponential divergence of $\xi$: $\xi\sim\exp(b/\sqrt{t})$, which is the signature of a BKT phase transition.
As the system approaches a BKT transition from a disordered phase, $\xi$ and $\chi$ diverge exponentially as $\sim\exp(b/\sqrt{t})$ [53, 54], where $t=(T-T_{c})/T_{c}$ is the reduced temperature, and $b$ is a non-universal constant. As shown in Fig. 11(a), a semi-logarithmic plot of $\xi$ versus $a+b\sqrt{t}$ is presented for $\sigma=2.1$ (purple), $3$ (red), and the NN case (black), where the non-universal constant $a$ comes from the amplitude of the scaling $\sim\exp(b/\sqrt{t})$, and the specific values of $a$ and $b$ are provided in the corresponding caption.
For a given $t$ that is not so close to the critical point, the growth of $\xi$ as a function of $t$ then reveals the thermodynamic growth law of correlation length. However, when the system enters the finite-size critical window, where $\xi\sim L$, the $\xi$ curve begins to bend to a plateau due to finite-size scaling behavior, thus deviating from the thermodynamic growth law. (Note that this analysis is different from the FSS analysis. Here, one first takes the $L\rightarrow\infty$ limit for a given temperature $T$ and then takes the $T\rightarrow T_{c}$ limit to examine the asymptotic growth law of $\xi$. In contrast, the FSS analysis first takes the $T\rightarrow T_{c}$ limit for a given $L$, followed by the $L\rightarrow\infty$ limit, studying the scaling behavior within the finite-size critical window.) As shown in Fig. 11(a), the straight black line in the semi-log plot clearly demonstrates the exponential growth of $\xi$ for $\sigma>2$, the typical BKT scaling behavior. In the large $t$ regime (the bottom left corner of Fig. 11(a)), where the system is far from criticality, $\xi$ slightly deviates from the exponential growth. Moreover, in Fig. 11(b), we further plot the ratio $\xi/L$ versus $b/\sqrt{t_{L}}=b/\sqrt{t(\ln L/L_{0})^{2}}$ to examine the FSS of the expoential growth, where $L_{0}$ is non-universal length scale, simply set as $1$. The scaling field $t_{L}=t[\ln(L/L_{0})]^{2}$ [55, 56] originates from the exponential divergence of the correlation length near the BKT transition $\xi\sim L\sim\exp(b/\sqrt{t})$. The data points collapse onto a single curve, showing that for $\sigma>2$, the phase transitions belong to the BKT universality class. For $\sigma=2.1$, taking into account that it is so close to the crossover point $\sigma=2$, it is somewhat surprising that its $\xi$ data can collapse well onto the curve for $\sigma=3$ and NN.
@@
Figure 15: The zoom-in of the second-moment correlation length ratio, $\xi/L$, as a function of inverse temperature $\beta$ for various system sizes at $\sigma$: $1.75$ (a), $1.875$ (b), and $2$ (c). Clear crossing behavior can be observed in all the cases, despite the shift of the crossing point due to the finite-size effect.
In Fig. 12(a), a semi-logarithmic plot of $a\xi$ versus $b/\sqrt{t}$ for various $\sigma$ is presented. The non-universal constants $a$ and $b$, obtained through the fitting, enhance the overall readability of plots without affecting the universal growth law. The specific values of $a$ and $b$ are provided in the corresponding caption. For $\sigma=2.1$ (purple), $\sigma=3$ (red), and the NN case (black), the straight black line demonstrates that the correlation length diverges exponentially $\xi\sim\exp(b/\sqrt{t})$, characterizing a typical BKT transition. For clarity, we exclude certain data points in the finite-size critical window, which manifests as a plateau. For $\sigma=1.875$ (green) and $2$ (blue), when sufficiently away from $\beta_{c}$, i.e. $b/\sqrt{t}$ is small, $\xi$ behaves seemingly like that for $\sigma=2.1$, $3$ and the NN case. However, near the critical point, i.e., as $b/\sqrt{t}$ approaches infinity, the behavior of $\xi$ becomes increasingly different from SR cases ($\sigma>2$) and deviates from the exponential growth, suggesting a different universality class. In fact, the growth of $\xi$ for $\sigma=1.875$ and $2$ can be approximately described by a power-law function (the curved lines in dark green and dark blue), indicating a second-order transition. In Fig. 12(b), a similar analysis is applied to the susceptibility, $\chi$. The plot shows $a\chi$ versus $b/\sqrt{t}$ on a semi-logarithmic scale. For the BKT cases ($\sigma=2.1$, $3$, and NN), $\chi$ diverges exponentially, with the straight black line guiding the eye. On the other hand, for $\sigma=1.875$ and $2$, $\chi$ follows a power-law growth near the critical point, with the dark-blue and dark-green curved lines guiding the eye, suggesting a second-order phase transition.
Figure. 13 presents the log-log plot of $a\xi$ and $a\chi$ versus ${b^{2}}/{t}$, where the x-axis is the square of ${b}/{\sqrt{t}}$. In Fig. 13(a), for $\sigma=1.875$ and $2$, the growth of $\xi$ asymptotically follows a linear relation as ${b^{2}}/{t}\rightarrow\infty$ (the straight dark-green and dark-blue lines), consistent with the power-law growth behavior of a second-order transition. The deviation of $\xi$ from linearity and the discrepancy between the exponent estimated from Fig. 13(a) and the fitting result in Table 5 are likely due to the strong correction to scaling from the irrelevant fields. Specifically, $\xi\sim t^{-\nu}g(u/t^{y_{u}/y_{t}})$, where $u$ is some irrelevant field and $y_{u}<0$ is the correpsonding scaling exponent. Nevertheless, the growth behavior of $\xi$ is already capable of clearly revealing the type of phase transition for various $\sigma$. For the growth of $\chi$ in Fig. 13(b), the straight dark-green and dark-blue lines illustrate a power-law behavior, as $\chi\sim t^{-\gamma}$ for $\sigma=1.875$ and $2$, again indicating a second-order transition.
The growth behavior of $\xi$ for $\sigma=2,1.875$ slightly deviates from a power law in Fig. 13(a), primarily because of the strong correction from the irrelevant fields. To mitigate such corrections, one needs to observe a larger $\xi$ closer to the critical point. However, brute-force simulations at larger systems are numerically challenging and impractical. Instead, one can exploit the FSS behavior of $\xi$ and extrapolate $\xi$ to obtain its thermodynamic value. The extrapolation method has been successfully employed to the classical 2D Heisenberg model [57, 58, 59, 44]. This method is based on the FSS ansatz:
$\frac{\mathcal{\xi}(\beta,sL)}{\mathcal{\xi}(\beta,~{}L)}=F_{\mathcal{\xi}}\left(\frac{\xi(\beta,L)}{L};s\right)+O(\xi^{-\omega},L^{-\omega})$ (38)
where $s$ is a fixed rescaling factor, typically set at $2$ for the convenience of MC simulations; $F_{\xi}$ is a universal function of $\xi/L$, depending on the rescaling factor $s$; the second term $O(\xi^{-\omega},L^{-\omega})$ contains further correction terms, which vanishes when $\xi$ and $L$ becomes large enough. In the ansatz (38), $F_{\xi}(x)$ typically takes the form:
$F_{\mathcal{\xi}}(x)=1+a_{1}e^{-1/x}+a_{2}e^{-2/x}+\cdots+a_{n}e^{-n/x}.$ (39)
with $n$ usually no more than $12$ [44]. The correction term $O(\xi^{-\omega},L^{-\omega})$ is negligible in the 2D Heisenberg case, which is also the case for the 2D LRXY model at $\sigma=3$. However, for $\sigma=2$, because of large finite-size correction, $O(\xi^{-\omega},L^{-\omega})$ term in ansatz (38) can not be ignored. The fitting equation for ansatz (38) is thus adjusted to:
$\frac{\mathcal{\xi}(\beta,sL)}{\mathcal{\xi}(\beta,~{}L)}=1+a_{1}e^{-1/x}+a_{2}e^{-2/x}+\cdots+a_{n}e^{-n/x}+bL^{-\omega}.$ (40)
The fitting results for $\sigma=2$ and $3$ are presented in Table 8 and 9 in Appendix 9. With the concrete form of Eq. (38), one can extrapolate the truncated $\xi$ to its thermodynamic value. See Ref. [44] for details of the method. The final extrapolation results, along with original data in Fig. 12 and 13, are shown in Fig. 14. The extrapolated $\xi$ reveals the characteristic signatures of the phase transition more clearly. In Fig. 14(a), $\xi$ is plotted vs. $1/t$ in the double-log scale. The linear behavior of the extrapolated data for $\sigma=2$ demonstrates a clear power-law divergence of $\xi$ with $\nu\approx 2.24$, characteristic of a second-order phase transition. In Fig. 14(b), $\xi$ is plotted vs. $1/\sqrt{t}$ in the semi-log scale, and the linearity of the extrapolated data for $\sigma=3$ demonstrates the exponential divergence of $\xi$, consistent with a BKT phase transition. Moreover, the extrapolated correlation lengths agree well with the original simulation data for both the $\sigma=2$ and $3$ cases, affirming the robustness and reliability of our extrapolation procedure.
In all, our analysis of the growth behavior of $\xi$ and $\chi$ strongly suggests that the system undergoes a second-order transition for $\sigma\leq 2$, while a BKT transition for $\sigma>2$; hence, the threshold between SR and LR universality is at $\sigma_{*}=2$.
VI Critical properties
@@
Figure 16: The extrapolation of $\xi/L$ curves’ crossing point $\beta_{L}$ of different system sizes. The $x$-axis represents the correction term $L^{-\omega}$: $L^{-0.86}$ in panel (a), $L^{-0.84}$ in panel (b), and $L^{-0.75}$ in panel (c); the $y$-axis denotes the inverse temperature $\beta$ at which the correlation length ratios for two consecutive system sizes intersect. The red line indicates the extrapolated value of $\beta$ as the system size approaches infinity.
@@
In the previous sections, we have studied the low-T and high-T properties of the system for $\sigma\leq 2$. In this section, we focus on the system’s critical properties for $\sigma\leq 2$. The fitting of the critical point and critical exponent is elaborated in subsection VI.1 and VI.2, respectively. Then, to determine the reliability of the estimate of $\eta$, the correlation function is also studied in subsection VI.3. Finally, we take a brief look at the properties of the specific-heat-like quantity in subsection VI.4, which provides additional hints of the crossover by showing distinct behaviors in the LR and SR universality.
VI.1 The Fit of Critical Point
@@
In this subsection, we focus on determining the critical temperature for various values of $\sigma$. For $\sigma\leq 2$, a second-order phase transition from a disordered phase to a ferromagnetic phase happens, where the dimensionless ratio curves, such as $Q_{m}$ and $\xi/L$, for different system sizes, should intersect at the critical point. Figure 15 presents $\xi/L$ as a function of inverse temperature $\beta$ for system sizes ranging from $L=64$ to $L=4096$, and for $\sigma=1.75$, $1.875$, and $2.0$. Unlike Fig. 2, which covers a broader range, this figure zooms in on a small regime of $\beta$ to observe the crossing behavior of $\xi/L$ for different system sizes. However, Fig. 15 shows that the crossing points for different system sizes do not intersect at a single value of $\beta$. Instead, these points shift toward higher values of $\beta$ as the system size increases, a phenomenon attributed to finite-size effects. The distance between the crossing points decreases as the system size grows, and with larger system sizes, the crossing points eventually converge at a single value of $\beta$, marking the critical point. To confirm this behavior and accurately determine the critical point, a finite-size analysis is performed at the crossing point $\beta_{L}$ between system sizes of $L$ and $L/2$.
@@
Figure 17: Demonstration of the reliability of the fitting of $\chi$. The semi-logarithmic plots of $\chi L^{\eta-2}-bL^{-\omega}$ versus $L$ for $\sigma=1.75,1.875,$ and $2$ are shown (where $\omega=0.48,0.42,0.5$ respectively). The red dots represent $\eta$ at the fitted center value $\eta_{c}$, while the green and blue ones represent $\eta$ deviating from the center value by positive and negative three standard deviations, respectively. The red square approaches a horizontal line as $L\rightarrow\infty$, while others show evident growing and decreasing trend, indicating the precise estimate for $\eta$ that we obtain.
@@
Figure 18: Demonstration of the reliability of the fitting of $K$. The semi-logarithmic plots of $KL^{-y_{t}}-bL^{-\omega}$ versus $L$ for $\sigma=1.75,1.875,$ and $2$ are shown (where $\omega=0.25,0.25,0.2$ respectively). The red dots represent $y_{t}$ at the fitted center value, while the blue and green ones represent $y_{t}$ deviating from the center value by positive and negative three standard deviations, respectively. The red square approaches a horizontal line as $L\rightarrow\infty$, while others show evident growing and decreasing trend, indicating the precise estimate for $y_{t}$ that we obtain (note that $y_{t}=1/\nu$). It is worth noting that, compared to $\chi$, since $K$ is defined by the ratio of two vanishing covariance and variance in Eq. (10), it exhibits larger errors. This is particularly evident for $L=8192$, where the difficulty in simulating longer Markov chain lengths results in significant errors for $K$ under this parameter. Therefore, the data point for $L=8192$ is not included in the fitting of $K$.
@@
where $\omega$ is related to the correlation length exponent $\nu$ and the leading irrelevant exponent in the renormalization group framework. Figure 16 illustrates the fitting results, with the $x$-axis representing $L^{-\omega}$ and the $y$-axis showing the corresponding values of $\beta_{L}$. In each case, the fitting lines exhibit finite intercepts on the $y$-axis, highlighted by red stripes. This confirms that the $\xi/L$ curves finally intersect at a single critical point as the system size increases. The intercepts provide an estimate of the critical temperature $\beta_{c}$. For further accuracy, the dimensionless ratios, such as $\xi/L$ and $Q_{m}$, are also fitted with respect to $L$ and $\epsilon=\beta-\beta_{c}$ near the critical point to refine the estimate of the critical point. Considering the universal part of dimensionless ratio $Q$ as in Eq. (16) and (17), Certain Taylor expansion provides a fitting equation of $Q$:
$Q(\epsilon,u,L)$ $=a_{0}+\sum_{i=1}^{m}a_{i}(\epsilon L^{y_{t}})^{i}+b_{1}L^{-\omega}+b_{2}L^{-2\omega}$ (42)
@@
where the value of $m$ depends on the specific fitting process and is typically taken as $2$ or $3$; $L^{-\omega}$ term comes from the irrelevant field $u$, and $\omega=-y_{u}$. In the formula above, terms such as $\epsilon^{2}L^{y_{t}},\epsilon^{3}L^{y_{t}},\epsilon^{3}L^{2y_{t}},\cdots$, are not under consideration since including such terms in the specific fitting process has almost no impact on the fit, and their amplitudes are also very close to zero. For the fit of $Q$, when $\sigma\leq 1.5$, $Q_{m}$ is more suitable for fitting, whereas when $1.5<\sigma\leq 2$, $\xi/L$ provides a better fit. By combining the two quantities, relatively accurate estimates of critical points can be obtained, and they are consistent with those derived from the fit of $\beta_{L}$. The final estimated values of $\beta_{c}$ are presented in Table 4 (Note that, except for $\sigma=1.25,1.5,1.75,1.875$ and $2$, the critical points under other parameters are only determined with much less precision since they are mainly used for drawing the phase diagram in Fig. 1).
Table 4: Estimations of critical point $\beta_{c}$ for various $\sigma$ shown in Fig. 1. Because of the normalization procedure in Eq. (3), when $\sigma\to-2$, $\beta_{c}\to\frac{1}{2}$, the critical point of XY model on the complete graph; when $\sigma\to\infty$, $\beta_{c}$ approaches the value of the NN case.
@@
When $\sigma>2$, the system undergoes a BKT transition, where the susceptibility $\chi$ is known to follow the scaling behavior: $\chi\sim L^{7/4}(\ln L+C_{1})^{1/8}$ at the critical point [60] where $C_{1}$ is a nonuniversal constant and needs to be obtained through fitting. Therefore, plotting $\chi L^{-7/4}(\ln L+C_{1})^{-1/8}$ as a function of temperature, the curves of different system sizes intersect at the critical point.
VI.2 The Fit of Critical Exponents
According to the scaling relationship, only two critical exponents are independent, and here we mainly focus on $\eta$ and the thermal scaling exponent $y_{t}=1/\nu$, with $\nu$ the correlation-length exponent. In Sak’s criterion, $\eta=\max(2-\sigma,\eta_{\rm{SR}})$ where $\eta_{\rm{SR}}=\frac{1}{4}$ for 2D XY model, so estimating the value of $\eta$ is another way to determine whether $\sigma_{*}$ equals 1.75 or 2. By fitting the susceptibility $\chi$ and the scaled covariance $K$, the values of $\eta$ and $y_{t}$ can be numerically estimated. At the critical point, considering the finite-size scaling Eq. (13) and (15), $\chi$ and $K$ are respectively fitted to:
$\chi=L^{2-\eta}(a+bL^{-\omega})+c$ (43)